The Math Club raised $4000 to attend a future competition. The club invested part of the money in an account that paid 5% interest, and the rest in a government savings bond that paid 4% interest per year. After one year, the club had earned a total of $170 in interest. How much was invested at each rate?
Answer by ankor@dixie-net.com(22740)

The Math Club raised $4000 to attend a future competition.
a + b = 4000
The club invested part of the money in an account that paid 5% interest, and the rest in a government savings bond that paid 4% interest per year.
After one year, the club had earned a total of $170 in interest.
.05a + .04b = 170
:
use elimination, multiply the above equation by 20, subtract from the 1st eq
a + b = 4000
a + .8b = 3400
--------------------subtraction eliminates a, find b
0 + .2b = 600
b = 600/.2
b = $3000 invested in b
How much was invested at each rate?
4000 - 3000 = $1000 invested in a
:
:
Check
.05(1000) + .04(3000) =
50 + 120 = 170
: